O_{2}(^{5}Π_{g}), ION-PAIR STATES, AND OXYGEN ATOM RECOMBINATION

\underline{D. L. HUESTIS}, C. G. BRESSLER, AND R. A. COPELAND, Molecular Physics Laboratory, SRI International, Menlo Park, CA 94025.

Collisions of oxygen or nitrogen molecules with laser-excitated high vibrational levels of O_{2}(A^{3}Σ_{u}^{+}) produce a longer lived excited state, whose resonant multiphoton ionization (REMPI) spectrum cannot be assigned to any known singlet or triplet system [1]. We attribute the lower state of this new transition to the predicted [2,3] but previously unobserved 5Π_{g} valence state of O_{2}. The regular sequence of vibrational bands observed suggests that the upper state is an ion-pair state of O_{2}, dissociating to O^{+} + O^{−}. Two vibrational levels have been observed in the lower state and ten in the upper state, the latter in the range 97,000 to 100,000 cm^{−1}.

Ab initio calculations [2,3] indicate that, although quite weakly bound, at large internuclear distances the ^{5}Π_{g} valence state should be the lowest of the states dissociating to ground state atoms. Because of its high degeneracy it could be a key intermediate state in O + O recombination [4,5]. We are developing a more detailed and quantitative theoretical model of O + O + M collisions. In addition to the experimental spectroscopy and kinetics results, we will also present preliminary work on spin-orbit resolved long-range potential curves for O_{2} and implications for oxygen atom recombination and the role of O_{2}(^{5}Π_{g}) in predicting atmospheric nightglow emissions [6].

References:

[1] C. G. Bressler and R. A. Copeland, “Collisional Production and Removal of O_{2} in the ^{5}Π_{g} State,” in preparation.

[2] R. P. Saxon and B. Liu, J. Chem. Phys. 67, 5432 (1977).

[3] H. Partridge, C. W. Bauschlicher, S. R. Langhoff, and P. R. Taylor, J. Chem. Phys. 95, 8292 (1991).

[4] P. C. Wraight, Planet. Space Sci. 30, 251 (1982).

[5] D. R. Bates, Planet. Space Sci. 36, 875 (1988).

[6] T. G. Slanger and D. L. Huestis, J. Geophys. Res. 86, 3551 (1981). Acknowledgments: This work is supported by the NASA Sun-Earth Connection and Planetary Atmospheres Programs.